OpenStudy (anonymous):

Solve this equation by FACTORING 16x(x-2)=8x-25 I would really like some help on figuring this out. Please and thank you!

OpenStudy (anonymous):

gunna work on it, give me a min or so.

OpenStudy (anonymous):

okay thank you so much!

OpenStudy (anonymous):

So at first you need to expand the left side and move everything there to have one side = 0 in order to factor. That being so, i got 16x^2 -32x - 8x +25 = 0 then collect all like terms 16x^2 -40x +25 = 0 I notice both the first and last numbers are squares and see that this is going to be a squared equation. So I factor from here to get (4x-5)^2

OpenStudy (anonymous):

Is the -40x included also?

OpenStudy (anonymous):

Hmm? In the factor? Yes. See (4x-5)^2 = (4x-5)(4x-5) Expanded u get 16x^2 -40x +25

OpenStudy (anonymous):

okay I see what you did. Thank you so much!

OpenStudy (anonymous):

Glad to help! :)